When Alex Roanguez moved to the Texas Rangers In 2001, ne received a lat of attentioN tor his $252 million contract (the to2011 2012 2013 2014 2015 2016 2017 2018 2019 2020 Deferred $3 M S3 M S3 M S3 M S5 M $4 M $3 M $3 M $3 M $3 M Assume that an aa. Calculate the true promised payments under this contract, including the deferred payments with interest values to the nearThe true promised payments under this contract for years 6 through 10 are: (Round all values to the nearest million.) Year 6:Year 11: $ million. Year 12: Smillion. Years 13 to 20: Smillion. b. Draw a timeline of all of the payments. The timeline of aO A. Year 2 8 10 11 12 13 20 4 - Cash Flows $18 $6.72 $5.38 $4.03 $19 $19 S19 S21 $19 $23 $23 $27 $27 S4.03 O B. Year 7 9 10OD. Year 1 6 7 8 9 10 11 12 13 20 - - Cash Flows $18 S4.03 $19 $19 $19 $21 $19 S23 $27 $27 S27 $6.72 $5.38 $4.03 c. CalculateThe reason for the difference is although the $252 million quoted value does discount the future cash flows, it does not adju
Answer #1

All figures in $ Mill ion Contract Signing Deferred Present Total Total Pay PVF(7%) Value Year Pay Bonus Pay Interest 16.8228

Note - 1

To calculate interest on deferred payments = For example for Year 2015(deferred from 2005 = 10 years)

=3(1+3%)10 -3 = 3(1.03)10-3 = 1.0317

Note - 2

To calculate Present Value Factor = 1/(1+rate)n

For example for Year 2001 =1/(1+0.07)1 =0.9346

For example for Year 2015 =1/(1+0.07)15 =0.3625

Answers:

A)18 million 19 million 19 million 19 million 21 million 19 million 23 million 27 million 27 million 27 million 6.7196 million

B)D. Year 2 34 5 6 7 8 10 11 12 13 20 Cash Flows $18 $19 $19 $19 $21 $19 $23 $27 $27 $27 $6.72 $5.38 $4.03 $4.03

C) Present Value of the contract = $165.7904 million

D) Option B. Because neither the interest on deferred payment was calculated nor the discounting is been done.